3-BUTENESELENOL: MICROWAVE SPECTRUM AND AB INITIO CALCULATIONS

D. PETITPREZ,\underline{G. WLODARCZAK}, J. DEMAISON, Laboratoire PhLAM, UMR CNRS 8523, Université de Lille 1, F-59655 Villeneuve d’Ascq, France; J.-C. GUILLEMIN, Laboratoire de Synthèse et Activation de Biomolécules, UMR CNRS 6052, ENSCR, F-35700, Rennes, France; H. MØLLENDAL, Department of Chemistry, The University of Oslo, NO-0315 Oslo, Norway.

3-Buteneselenol H_{2}C=CHCH_{2}CH_{2}SeH was investigated by microwave Fourier transform (5-20 GHz) and microwave Stark (24-62 GHz) spectroscopies. The spectra look rather complex due to the large number of possible conformers (15) and the number of isotopic species for selenium (6). Moreover in the Stark spectra the low-lying excited states are also detected. Quantum mechanical calculations (G2 method at the MP2 and B3LYP levels of theory) were carried to precise the relative energies of all these conformers. Their geometrical structures, centrifugal distortion constants and dipole moments were also calculated, leading to the identification of the three more stable conformers. The most stable conformer exhibits an intramolecular hydrogen bond involving the selenium atom.